the first of 3 numbers is 3 times the second and the third is 8 less than 2 times the second. their sum is 130. find the numbers?
Answer by checkley77(12569)

x+y+z=130
x=3y
z=2y-8
3y+y+(2y-8)=130
6y=130+8
6y=138
y=138/6
y=23 ans for the second number.
x=3*23=69 for the first number
z=2*23-8=46-8=38 for the third number.
Proof:
23+69+38=130
130=130